What is perpendicular to y=4x-7 through (5,-1)
Answer by rodriguezh2(15)

When doing these problems always make sure your given equation is in y = mx + b form. Which yours is y = 4x - 7
Remember (m) represents slope and (b) represents the y-intercept.
Perpendicular lines have negative reciprocal slopes. Your slope for your given line is 4. This means the line we are looking for that is perpendicular to the given line is -(1/4)
We also know that your line contains the point (5,-1)
Remember that 5 represents the x value and -1 represents the y value
We can plug these numbers into the equation: (y - y1) = m(x - x1)
(y - (-)1) = -(1/4)(x - 5) Remember to change your -1 to a positive number
(y + 1) = -(1/4)x + (5/4) Multiply the right side of the equation through
Remember that when you multiply two negatives you get a positive number
Now subtract 1 from both sides; to do this you need to change one to (4/4) getting the least common denominator so you can subtract: (5/4) - (4/4)
So we are left with:
y = -(1/4)x + 1/4
Hope this helps!